# Causal Data Fusion for Panel Data without a Pre-Intervention Period

### v3

@@ 5 Simulation Studies
For the linear equi-confounding approach, let $F_{0}=\frac{1}{J}\sum_{i=2}^{J+1}F_{i}$ and $Y_{0}=\frac{1}{J}\sum_{i=2}^{J+1}Y_{i}$. Note that Assumption 2 need not hold for the factor model in Equation (10). For instance, Figure 1(a) shows one realization of the DGP ($T=20$) in which the line connecting $\mathbb{E}[F_{1}]$ to $\mathbb{E}[Y_{1}^{(0)}]$ is not parallel to the line connecting $\mathbb{E}[F_{0}]$ to $\mathbb{E}[Y_{0}]$. Hence, the estimator $\hat{\psi}^{{eq}1}$ is expected to be biased. However, across extensive simulations (multiple DGP realizations), we observed that with our choice of distributions and parameters, the bias normalized by the ground-truth causal effect was negligible in most cases. A similar concern holds for the logarithmic equi-confounding approach: Assumption 3 need not hold. Specifically, the lines connecting $\log\{\mathbb{E}[F_{1}]\}$ to $\log\{\mathbb{E}[Y_{1}^{(0)}]\}$ and $\log\{\mathbb{E}[F_{0}]\}$ to $\log\{\mathbb{E}[Y_{0}]\}$ are not parallel (see Figure 1(b)). Nevertheless, comparing Figures 1(a) and 1(b), we see a smaller violation of the logarithmic equi-confounding assumption in the corresponding DGP.
@@
Turning to the synthetic control method, note that the optimization problem in Equation (9) has a quadratic objective with quadratic constraints. We solve it using the gurobi package (Gurobi Optimization, LLC,, 2024) in R. A key aspect of our methodology is the trade-off between the matching quality in the reference and target domains. This balance is governed by the flexibility allowed via the parameters $\eta_{Z}$ and $\eta_{X}$, which control the relative importance of matching in the reference and target domains, respectively. A larger $\eta_{Z}$ allows better matching in the target domain at the expense of the reference domain, whereas a larger $\eta_{X}$ allows better alignment in the reference domain but may reduce the quality of matching in the target domain. This trade-off can be tuned based on domain-specific considerations regarding the association between covariates and outcomes. In our implementation, we set both $\eta_{Z}$ and $\eta_{X}$ to $0.1$. We also conducted a sensitivity analysis to examine the dependence of the estimated causal effect on $\eta_{Z}$ and $\eta_{X}$. Figure 2(a) presents the results of this analysis. As seen in the figure, deviations from 0.10.1—either increasing or decreasing $\eta_{Z}$ and $\eta_{X}$—exhibit minimal impact on the estimated causal effects.
@@
The same DGP described above is used for the synthetic control method. We draw a single dataset with $T=20$. In Figure 3(a), the black solid line shows the trajectory of the observed target unit and the black dashed line shows the trajectory of $\sum_{i=2}^{J+1}w^{*}_{i}F_{i,t}$ for $t\in\{1,\ldots,T\}$. The gray solid lines represent the observed control units for $i\in\{2,\ldots,J+1\}$. We observe that the synthetic control target unit closely tracks the observed target unit in the reference domain. In Figure 3(b), the black and gray solid lines represent the paths of the observed units for $i\in\{1,\ldots,J+1\}$ and $s\in\{1,\ldots,S\}$, and the solid red line represents the counterfactual target unit in the target domain. The dashed line shows the synthetic control target unit $\sum_{i=2}^{J+1}w^{*}_{i}Y_{i,s}$ using the same weights $w^{*}$ as above. Figure 3(b) demonstrates that the synthetic control data fusion method provides a good approximation to $Y_{1,s}^{(0)}$ using weights $w^{*}$ learned from the reference domain.
@@
Figure 4: Bias and NSEs in synthetic control and equi-confounding methods.
Figure 4(c) provides a comparative visualization of the biases from the synthetic control data fusion and equi-confounding data fusion methods. We present the biases averaged over $300$ datasets for each value of $T$, along with the interquartile range. Overall, we observe superior performance from the synthetic control data fusion method. The bias from the synthetic control data fusion method decreases as $T$ increases, eventually approaching zero, reaffirming Theorem 4 that the bias vanishes as the number of reference-domain periods grows. The logarithmic equi-confounding method also exhibits a decreasing bias trend. However, there is no clear decreasing pattern for the bias from the linear equi-confounding method.
@@ 5.1 Placebo Test
To evaluate the ability of the synthetic control method to predict counterfactual outcomes, we implement a placebo test, an inferential technique originating in (Abadie and Gardeazabal,, 2003). Specifically, we apply the synthetic control method to each unit in the donor pool by generating synthetic controls using the same procedure employed for the target unit, shifting the target unit into the donor pool. We then compare the distribution of these placebo effects with the effect estimated for the target unit. By comparing the distribution of estimated placebo effects to the target-unit estimate, we can assess whether the observed effect is significantly larger than would be expected by chance.
@@ 6 Application to COVID-19 Vaccination Rate Analysis
Results: We used the value $0.1$ for both $\eta_{Z}$ and $\eta_{X}$. However, as in the simulation study, we performed sensitivity analysis to investigate the dependence of our estimated causal effect on the choice of $\eta_{Z}$ and $\eta_{X}$, and we observed a small impact of deviations from these hyperparameters. Figure 2(b) presents the sensitivity analysis. Table 1 presents the covariates used in the estimation process. As seen there, the method successfully generated the synthetic Chelsea with covariates close to those of the real Chelsea. Table 2 displays the weights assigned to each control city in constructing synthetic Chelsea. The results indicate that the best matching for Chelsea is achieved by combining Revere, Holyoke, Everett, Lawrence, and Springfield. All other units in the donor pool were assigned zero weights.
@@
Covariates Chelsea Synthetic Chelsea Average of 19 Controls
@@
Figure 6 presents the COVID-19 vaccination rates for the observed units and synthetic Chelsea in the reference and target domains. In Figure 6(a), synthetic Chelsea closely follows actual Chelsea in the reference domain, demonstrating that our algorithm has succeeded in obtaining a good match. This, along with the high degree of balance in observed covariates, supports the validity of synthetic Chelsea as a counterfactual for actual Chelsea in the absence of the intervention. In Figure 6(b), the observed Chelsea Hispanic COVID-19 vaccination rate and its synthetic version begin to diverge after the intervention starts, with the gap widening over time. Specifically, Chelsea exhibits a steeper rise in the earlier months. The divergence between the black solid line (actual Chelsea) and the black dashed line (synthetic Chelsea) suggests that La Colaborativa’s efforts contributed to a greater increase in Hispanic vaccination rates than would have been expected without the intervention. We also provide the results of the placebo test in Figures 5(c) and 5(d), which confirm the statistical significance of our findings regarding the causal effect.
We now present the estimated causal effect of the community organization intervention based on our proposed methodologies. Over the 16-month period from March 2021 to July 2022, the COVID-19 vaccination rate for the Hispanic sub-population is estimated to have increased by $13.6\%$ using the linear equi-confounding method, $13.2\%$ using the logarithmic equi-confounding method, and $10.1\%$ using the synthetic control data fusion method. All three methods indicate a similarly large positive causal effect, affirming the effectiveness of the community organization intervention.
@@ A.1 Proof of Theorem 1
A.2 Proof of Theorem 2
From Theorem 3.2, the parameter of interest $\psi_{0}$ is identified as
@@ A.2 Proof of Theorem 2
Hence, using Parts (1) and (3) of Assumption 4, the absolute bias of $\hat{\psi}^{eq2}$ will be bounded as
@@
A.3 Proof of Corollary 2
The proof is the same as that of Theorem 3.3, except for our treatment of $\left|cov\left(A,\frac{1}{B}\right)\right|$ in Display (12), for which we argue as follows:
@@ A.3 Proof of Corollary 2
A.4 Proof of Theorem 4
Consider the model presented in Assumption 5:
@@ A.4 Proof of Theorem 4
Note that under the above model, the potential outcome $Y_{i,s}$ under control will be
@@
The weighted average of the potential outcomes of the donors is
@@
$Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}=\varphi_{s}^{\top}(X_{1}-\sum_{i=2}^{J+1}w_{i}X_{i})+\vartheta_{s}^{\top}(\mu_{1}-\sum_{i=2}^{J+1}w_{i}\mathbf{\mu}_{i})+\tilde{\vartheta}_{s}^{\top}(\tilde{\mu}^{Y}_{1}-\sum_{i=2}^{J+1}w_{i}\tilde{\mu}^{Y}_{i})+\varepsilon_{1,s}-\sum_{i=2}^{J+1}w_{i}\varepsilon_{i,s}.$ (16)
Let $F_{i}^{R}$ be the $T\times 1$ vector with the $t$-th element equal to $F_{i,t}$. Similarly, let $\epsilon_{i}^{R}$ be the $T\times 1$ vector with the $t$-th element equal to $\epsilon_{i,t}$. Finally, let $\phi^{R}$, $\theta^{R}$, and $\tilde{\theta}^{R}$ be the $T\times d_{r}$, $T\times d_{u}$, and $T\times d^{F}_{\tilde{u}}$ matrices with the $t$-th rows equal to $\phi_{t}$, $\theta_{t}$, and $\tilde{\theta}_{t}$, respectively. We have
$F_{1}^{R}-\sum_{i=2}^{J+1}w_{i}F_{i}^{R}=\phi^{R}(Z_{1}-\sum_{i=2}^{J+1}w_{i}Z_{i})+\theta^{R}(\mu_{1}-\sum_{i=2}^{J+1}w_{i}\mu_{i})+\tilde{\theta}^{R}(\tilde{\mu}^{F}_{1}-\sum_{i=2}^{J+1}w_{i}\tilde{\mu}^{F}_{i})+\epsilon_{1}^{R}-\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R}.$
@@
$\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{F_{1}^{R}-\sum_{i=2}^{J+1}w_{i}F_{i}^{R}\}=$ $~\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\phi^{R}(Z_{1}-\sum_{i=2}^{J+1}w_{i}Z_{i})$
@@
$\vartheta_{s}^{\top}(\mu_{1}-\sum_{i=2}^{J+1}w_{i}\mu_{i})=$ $~\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{F_{1}^{R}-\sum_{i=2}^{J+1}w_{i}F_{i}^{R}\}$ (17)

### v4

@@ 5 Simulation Studies
For the linear equi-confounding approach, let $F_{0}=\frac{1}{J}\sum_{i=2}^{J+1}F_{i}$ and $Y_{0}=\frac{1}{J}\sum_{i=2}^{J+1}Y_{i}$. By choosing the parameters $\varphi_{s}$ and $\phi_{t}$ to be 0 for all $s$ and $t$, and requiring $\frac{1}{T}\sum_{t=1}^{T}\theta_{t}$ to be equal to $\frac{1}{S}\sum_{s=1}^{S}\vartheta_{s}$, Assumption 2 is satisfied: as shown in Figure 1(a) (which is based on a realization of the DGP with $T=20$), the line connecting $\mathbb{E}[F_{1}]$ to $\mathbb{E}[Y_{1}^{(0)}]$ is parallel to the line connecting $\mathbb{E}[F_{0}]$ to $\mathbb{E}[Y_{0}]$. By choosing the parameters $\varphi_{s}$ and $\phi_{t}$ to be 0 for all $s$ and $t$, and requiring $(\frac{1}{T}\sum_{t=1}^{T}\rho_{t})/(\frac{1}{S}\sum_{s=1}^{S}\varrho_{s})$ to be equal to $(\frac{1}{T}\sum_{t=1}^{T}\theta_{t})/(\frac{1}{S}\sum_{s=1}^{S}\vartheta_{s})$, Assumption 3 is satisfied: as shown in Figure 1(b) (which is based on the realization of the DGP with $T=20$), the lines connecting $\log\{\mathbb{E}[F_{1}]\}$ to $\log\{\mathbb{E}[Y_{1}^{(0)}]\}$ and $\log\{\mathbb{E}[F_{0}]\}$ to $\log\{\mathbb{E}[Y_{0}]\}$ are parallel.
Note that Assumptions 2 and 3 need not hold for the factor model in Equation (10) in general. To assess the robustness of the equi-confounding estimators, we evaluated the degree to which Assumptions 2 and 3 are violated under the factor model specified in Assumption 5 with the aforementioned choices of the distributions. We considered 1000 DGPs, holding $\mu_{i}$, $Z_{i}$, $X_{i}$ fixed while varying time-specific parameters $\rho_{t}$, $\phi_{t}$, $\theta_{t}$, $\varrho_{s}$, $\varphi_{s}$, $\vartheta_{s}$ and errors across DGPs. The empirical distribution of linear equi-confounding violations exhibits a mean of 0.871 (SD = 1.100), while logarithmic equi-confounding violations show substantially smaller deviations with a mean of 0.014 (SD = 0.017). Both distributions are centered near zero, indicating that the assumptions hold approximately on average. However, the logarithmic equi-confounding assumption demonstrates markedly tighter concentration around zero, suggesting it is substantially less sensitive to the particular realization of the DGP parameters.
@@
Turning to the synthetic control method, note that the optimization problem in Equation (9) has a quadratic objective with quadratic constraints. We solve it using the gurobi package (Gurobi Optimization, LLC, 2024) in R. A key aspect of our methodology is the trade-off between the matching quality in the reference and target domains. This balance is governed by the flexibility allowed via the parameters $\eta_{Z}$ and $\eta_{X}$, which control the relative importance of matching in the reference and target domains, respectively. A larger $\eta_{Z}$ allows better matching in the target domain at the expense of the reference domain, whereas a larger $\eta_{X}$ allows better alignment in the reference domain but may reduce the quality of matching in the target domain. This trade-off can be tuned based on domain-specific considerations regarding the association between covariates and outcomes. In our implementation, we set both $\eta_{Z}$ and $\eta_{X}$ to $0.1$. We also conducted a sensitivity analysis to examine the dependence of the estimated causal effect on $\eta_{Z}$ and $\eta_{X}$. As seen in the Figure 2(a), deviations of $\eta_{Z}$ and $\eta_{X}$ from 0.10.1–either increasing or decreasing–exhibit minimal impact on the estimated causal effects. Specifically, they lead to changes in the estimated causal effects that are bounded by $0.005$.
@@
We draw a single dataset with $T=20$. In Figure 3(a), the black solid line shows the trajectory of the observed target unit and the black dashed line shows the trajectory of $\sum_{i=2}^{J+1}w^{*}_{i}F_{i,t}$ for $t\in\{1,\ldots,T\}$. The gray solid lines represent the observed control units for $i\in\{2,\ldots,J+1\}$. We observe that the synthetic control target unit closely tracks the observed target unit in the reference domain. In Figure 3(b), the black and gray solid lines represent the paths of the observed units for $i\in\{1,\ldots,J+1\}$ and $s\in\{1,\ldots,S\}$, and the solid red line represents the counterfactual target unit in the target domain. The dashed line shows the synthetic control target unit $\sum_{i=2}^{J+1}w^{*}_{i}Y_{i,s}$ using the same weights $w^{*}$ as above. Figure 3(b) demonstrates that the synthetic control data fusion method provides a good approximation to $Y_{1,s}^{(0)}$ using weights $w^{*}$ learned from the reference domain.
@@
Figure 4: Bias and NSEs in synthetic control and equi-confounding methods (over 300 simulations) with a DGP satisfying Assumption 5.
(a) NSEs of $\mu_{i}$, $Z_{i}$, $X_{i}$, and $\bar{F}_{i}$.
(b) NSEs of $\mu_{i}$, $Z_{i}$, and $X_{i}$.
Figure 5: Bias and NSEs for synthetic control and equi-confounding methods (over 300 simulations) with a DGP satisfying the requirements of all frameworks.
Figure 4(c) provides a comparative visualization of the biases from the synthetic control data fusion and equi-confounding data fusion methods. We present the biases averaged over $300$ datasets for each value of $T$, along with the interquartile range. Overall, we observe superior performance from the synthetic control data fusion method. The bias from the synthetic control data fusion method decreases as $T$ increases, eventually approaching zero, reaffirming Theorem 4 that the bias vanishes as the number of reference-domain periods grows. The logarithmic equi-confounding method also exhibits a decreasing bias trend. However, there is no clear decreasing pattern for the bias from the linear equi-confounding method. This is due to the fact that the main assumption of this method is not necessarily satisfied. To examine this point, in Figure 5, we show performances of three proposed methods where the DGPs satisfy the requirements of all three methods.
@@
(a) Leave-one-out sensitivity test (reference domain).
(b) Leave-one-out sensitivity test (target domain).
(c) Leave-one-out sensitivity test for the COVID-19 vaccination rates application for Black sub-population (reference domain).
(d) Leave-one-out sensitivity test for the COVID-19 vaccination rates application for Hispanic sub-population (target domain).
Figure 7: Leave-one-out tests.
@@ 5.1 Placebo Test
To evaluate the ability of the synthetic control method to predict counterfactual outcomes, we implement a placebo test, an inferential technique originating in (Abadie and Gardeazabal, 2003). Specifically, we apply the synthetic control method to each unit in the donor pool by generating synthetic controls using the same procedure employed for the target unit, shifting the target unit into the donor pool. We then compare the distribution of these placebo effects with the effect estimated for the target unit. By comparing the distribution of estimated placebo effects to the target-unit estimate, we can assess whether the observed effect is significantly larger than would be expected by chance.
@@
We also conducted a leave-one-out placebo test by iteratively removing each donor unit with a positive weight ($w>0$) and re-estimating the synthetic control. Figures 7(a) and 7(b) demonstrate that the leave-one-out estimates closely track the original synthetic treated unit. This confirms the robustness of our results across both the reference and target domains for the simulated data.
@@ 6 Application to COVID-19 Vaccination Rate Analysis
Results: We used the value $0.1$ for both $\eta_{Z}$ and $\eta_{X}$. However, as in the simulation study, we performed sensitivity analysis to investigate the dependence of our estimated causal effect on the choice of $\eta_{Z}$ and $\eta_{X}$, and we observed a small impact of deviations from these hyperparameters. Figure 2(b) presents the sensitivity analysis. As seen in this figure, deviations of $\eta_{Z}$ and $\eta_{X}$ from 0.10.1–either increasing or decreasing–exhibit minimal impact on the estimated causal effects. Specifically, they lead to changes in the estimated causal effects that are bounded by $0.005$. Table 1 presents the covariates used in the estimation process. It shows that the proposed synthetic control method successfully identifies a combination of control units that closely approximates the covariates of Chelsea. For example, the synthetic Chelsea is significantly closer than the average of all control cities across different covariates. Although some discrepancies remain, such as in the “Proportion - Hispanic”. This emphasizes the importance of using a method that allows approximate matching as in our Equations (4)–(6). Table 2 displays the weights assigned to each control city in constructing synthetic Chelsea. The results indicate that the best matching for Chelsea is achieved by combining Revere, Holyoke, Everett, Lawrence, and Springfield. All other units in the donor pool were assigned zero weights.
@@
Covariates Chelsea Synthetic Chelsea Avg. of controls (n=19)
@@
Figure 8 presents the COVID-19 vaccination rates for the observed units and synthetic Chelsea in the reference and target domains. In Figure 8(a), synthetic Chelsea closely follows actual Chelsea in the reference domain, demonstrating that our algorithm has succeeded in obtaining a good match. This, along with the approximate matching in observed covariates, supports the validity of synthetic Chelsea as a counterfactual for actual Chelsea in the absence of the intervention. In Figure 8(b), the observed Chelsea Hispanic COVID-19 vaccination rate and its synthetic version begin to diverge after the intervention starts, with the gap widening over time. Specifically, Chelsea exhibits a steeper rise in the earlier months. The divergence between the black solid line (actual Chelsea) and the black dashed line (synthetic Chelsea) suggests that La Colaborativa’s efforts contributed to a greater increase in Hispanic vaccination rates than would have been expected without the intervention. We also provide the results of the placebo test in Figures 6(c) and 6(d), which confirm the statistical significance of our findings regarding the causal effect. We also conducted the leave-one-out placebo test, leaving out one of the five donor cities (Revere, Holyoke, Everett, Springfield and Lawrence) from the donor pool. In Figures 7(c) and 7(d), the results show the five additional synthetic controls closely track the original synthetic Chelsea in both reference and target domain.
The estimated causal effect of the community organization intervention based on our proposed methodologies is as follows. Over the 16-month period from March 2021 to July 2022, the COVID-19 vaccination rate for the Hispanic sub-population is estimated to have increased by $13.6\%$ using the linear equi-confounding method, $13.2\%$ using the logarithmic equi-confounding method, and $10.1\%$ using the synthetic control data fusion method. All three methods indicate a similarly large positive causal effect, affirming the effectiveness of the community organization intervention.
@@ Appendix A Proofs
Proof of Theorem 2
Due to the consistency assumption, $\psi_{0}$ can be written as
By Assumption 3,
$\mathbb{E}[Y_{1}^{(0)}]$ $=\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\sum_{i=2}^{J+1}F_{i}]}\mathbb{E}[\sum_{i=2}^{J+1}Y_{i}^{(0)}].$
Therefore, applying the consistency assumption one more time, we obtain
$\psi_{0}$ $=\mathbb{E}[Y_{1}]-\mathbb{E}[Y_{1}^{(0)}]$
$=\mathbb{E}[Y_{1}]-\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\sum_{i=2}^{J+1}F_{i}]}\mathbb{E}[\sum_{i=2}^{J+1}Y_{i}].$
Proof of Theorem 3
From Theorem 2, the parameter of interest $\psi_{0}$ is identified as
@@
Hence, using Assumption 4, the absolute bias of $\hat{\psi}^{eq2}$ will be bounded as
@@
Proof of Corollary 2
The proof is the same as that of Theorem 3, except for our treatment of $\left|cov\left(A,\frac{1}{B}\right)\right|$ in Display (12), for which we argue as follows:
@@
Proof of Theorem 4
Consider the factor models in Assumption 5:
@@
Under this model, the potential outcome under control satisfies
@@
Define $\bar{F}_{i}:=[F_{i,1}~\cdots~F_{i,T}]^{\top}$ and $\epsilon_{i}^{R}:=[\epsilon_{i,1}~\cdots~\epsilon_{i,T}]^{\top}$. Let $\phi^{R}$, $\theta^{R}$, and $\tilde{\theta}^{R}$ be the $T\times d_{r}$, $T\times d_{u}$, and $T\times d^{F}_{\tilde{u}}$ matrices whose $t$th rows are $\phi_{t}^{\top}$, $\theta_{t}^{\top}$, and $\tilde{\theta}_{t}^{\top}$, respectively.
Fix $s\in\{1,\ldots,S\}$. The weighted average of the potential outcomes of the donors is
@@
$Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}=\varphi_{s}^{\top}\Big(X_{1}-\sum_{i=2}^{J+1}w_{i}X_{i}\Big)+\vartheta_{s}^{\top}\Big(\mu_{1}-\sum_{i=2}^{J+1}w_{i}\mu_{i}\Big)+\tilde{\vartheta}_{s}^{\top}\Big(\tilde{\mu}_{1}^{Y}-\sum_{i=2}^{J+1}w_{i}\tilde{\mu}_{i}^{Y}\Big)+\varepsilon_{1,s}-\sum_{i=2}^{J+1}w_{i}\varepsilon_{i,s}.$ (16)
Moreover,
$\bar{F}_{1}-\sum_{i=2}^{J+1}w_{i}\bar{F}_{i}=\phi^{R}\Big(Z_{1}-\sum_{i=2}^{J+1}w_{i}Z_{i}\Big)+\theta^{R}\Big(\mu_{1}-\sum_{i=2}^{J+1}w_{i}\mu_{i}\Big)+\tilde{\theta}^{R}\Big(\tilde{\mu}_{1}^{F}-\sum_{i=2}^{J+1}w_{i}\tilde{\mu}_{i}^{F}\Big)+\epsilon_{1}^{R}-\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R}.$
@@
$\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{\bar{F}_{1}-\sum_{i=2}^{J+1}w_{i}\bar{F}_{i}\}=$ $~\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\phi^{R}(Z_{1}-\sum_{i=2}^{J+1}w_{i}Z_{i})$
@@
$\vartheta_{s}^{\top}(\mu_{1}-\sum_{i=2}^{J+1}w_{i}\mu_{i})=$ $~\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{\bar{F}_{1}-\sum_{i=2}^{J+1}w_{i}\bar{F}_{i}\}$ (17)